P7-5 Coupon Rates [LO2]

Kiss the Sky Enterprises has bonds on the market making annual payments, with 19 years to maturity, and selling for $840. At this price, the bonds yield 7.9 percent. What must the coupon rate be on the bonds?


rev: 09_18_2012

Multiple Choice

  • 12.49%

  • 6.25%

  • 7.44%

  • 7.90%

  • 6.35%
Answer #1

Current price=Annual coupon*Present value of annuity factor(7.9%,19)+$1000*Present value of discounting factor(7.9%,19)

840=Annual coupon*9.673082653+$1000*0.23582647

Annual coupon=(840-235.82647)/9.673082653

=62.46(Approx).

Hence coupon rate=Annual coupon/Face value

=62.46/$1000

=6.25%(Approx).

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=Annual coupon[1-(1.079)^-19]/0.079

=Annual coupon*9.673082653

2.Present value of discounting factor=1000/1.079^19

=1000*0.23582647